Find all the roots in C of the equation x^4-2x^3+4x^2+6x-21=0, give n that two of its roots are equal in magnitude and opposite in sign.
1
Expert's answer
2020-12-10T19:59:31-0500

There is a general solution to the 4th power equation, but it is very complex. Fortunately this equation can be solved quite easily.


"x^4-2x^3+4x^2+6x-21=(x^2+a)(x^2+bx+c)=0"


"x^4+bx^3+(a+c)x^2+abx+ac=0"


"b=-2, a+c=4,ab=6,ac=-21"


"b=-2, ab=6,a=-3,c=4-a=4+3=7"


Therefore the equation is:


"(x^2-3)(x^2-2x+7)=0"


So either


"x=\pm\sqrt{3}" or


"x^2-2x+7=0"


"D=b^2-4ac=4-28=-24,\sqrt{D}=2\sqrt{6}i"


"x=\frac{2\pm2\sqrt{6}i}{2}=1\pm\sqrt{6}i"


Out of 4 roots that we found "(\pm\sqrt{3},1\pm\sqrt{6}i)" only the first two are equal in magnitude and opposite in sign, so n=1 (one pair).